# Simultaneous Modeling of Protein Conformation and Dynamics via Autoregression

### v1

@@ Appendix B Diffusion Models on SE(3) Space
Diffusion Probabilistic Models (DPM) model complex distributions through iterative denoising. In the context of protein conformations, DPMs defined over $\mathrm{SE(3)}$ translation-rotation space have been applied for protein backbone structural generation [43, 40]. Following Section 2.2, $\mathbf{x}_{0}=(\mathbf{T}_{0},\mathbf{R}_{0})\in\mathrm{SE(3)}^{N}$ denotes the translations and rotations of backbone rigids in data. The diffusion processes defined in the translation and rotation subspace add noise to corrupt the data:
@@
[…]
@@ D.1 Detailed Module Architectures
To encode the structural information of each conformation frame, we introduced the FrameEncoder, a pseudo-beta-carbon (Cβ) coordinate encoder similar to the InputEmbedding module from AlphaFlow [17] (without diffusion time embedding). Specifically, this module first compute the pairwise distances between residues using Cβ coordinates. These distances are then binned, embedded into latent embedding, and further refined through triangular update blocks including triangle attention and multiplication updates [19]. See Algorithm 1 for the specifics. The resulting per-frame Cβ pair embedding $\mathbf{z}_{\text{Frame}}^{l}$ is concatenated with the pair representation from FoldingModule.
@@
In the Trajectory Module, we interleave layers of StructuralUpdate and TemporalUpdate to iteratively update the latent $[\mathbf{s}^{l},\mathbf{z}^{l}]$, enabling the temporal reasoning across frames and structural refinement within each frame.
For the StructuralUpdate, we adopt a Pairformer block from AlphaFold 3 [2], which jointly updates the single and pair embeddings of the current frame through structural reasoning. After the StructuralUpdate, the pair embedding is flattened from $[N,N,d]$ to $[N\times N,d]$ and concatenated with the single embedding before being passed into the TemporalUpdate.
To model temporal dependencies between frames, we use a lightweight Llama architecture [38]. We transpose the input such that the temporal dimension is treated as the sequence axis for channel-wise self-attention across time. Rotary positional encoding [37] is applied to encode the temporal position for each frame. A causal attention mask is applied to restrict each frame to only attend to previous frames. After the temporal update, the latent embedding are reshaped and split back into single and pair embeddings.
@@
Figure 9: Architecture details of the Trajectory Module. Trajectory Module contains interleaving blocks of StructuralUpdate and TemporalUpdate (only one block of each is shown). StructuralUpdate leverages the Pairformer architecture from Abramson et al. [1], updating the pair embeddings with triangular updates and the single embeddings using with pair bias from the updated pair. The updated pair embeddings are flatten and concatenate with single embedding for channel-wise temporal update. The attention is applied along the temporal dimension and update each single and pair embedding channels independently. The embeddings from TemporalUpdate are split and reshape back into single and pair embeddings.
@@
Conformation generation, conditioned on the temporal signals from the Trajectory Module, is performed using the model architecture from ConfDiff [40], As shown in Figure 10, following ConfDiff, the inputs to the denoising model include diffusion time $t$, pairwise distance between residue rigids, and residue indices (not shown). They are encoded and concatenated with the single and pair embeddings from FoldingModule. In addition, the single and pair embeddings from the Trajectory Module are projected back to the latent dimension $d_{\mathbf{s}}$ and $d_{\mathbf{z}}$, respectively, to modify the single and pair embeddings used by ConfDiff.
@@ D.2 Training and Inference details
We train all ConfRover models on the trajectories from the ATLAS training set, following the train-validation-test split of previous works [17, 40, 18]. Specifically, we exclude the training proteins longer than 384 amino acid residues, leading to 1080 training proteins.
@@ E.1 Trajectory Simulation: multi-start
Here we also include the scatterplot of Pearson correlation in Figure 11 and additional metrics from the multi-start experiments at different strides in Table 8. Across different strides, ConfRover models consistently outperforms MDGen in recovering the correct level of dynamics.
Figure 11: Scatterplots of conformation changes in sample trajectories versus those in the reference trajectories, measured by trajectory-level conformation changes, frame-level conformation changes, and next-frame difference ($\Delta$Frame), measured by the RMSD of alpha carbons (unit: Å) or $L^{2}$ distance in the projected PCA space. MDGen tends underestimate the magnitude of conformation changes while ConfRover generate samples at similar level as the MD reference. The exact match of measured conformation changes is not possible due to stochastic sampling in both MD simulation and generative models.
Table 8: Additional metrics from the multi-start benchmark. Results for different strides are shown in separate blocks. The better score in each block is highlighted in bold (excluding diversity)
@@ E.3 Time-independent Conformation Sampling
We follow the evaluation protocol of Ye et al. [42] to assess time-independent conformation sampling on the ATLAS test set. For each protein, 250 independent samples are generated. Since MDGen does not support time-independent sampling, we approximate its performance using samples from its 100-ns trajectory, serving as a sequential-sampling baseline. The performance of state-of-the-art models, AlphaFlow and ConfDiff, is taken from Table 10 of Ye et al. [42], using their best-performing variants: AlphaFlow-MD and ConfDiff-Open-MD. Full results are shown in Table 9.
Table 9: Performance on time-independent conformation generation on ATLAS. A total of 250 conformations were sampled for each protein, and the median values across 82 proteins are reported. The best performance is highlighted in bold, and the second-best is underlined.*MDGen does not support time-independent sampling and the result from sequential sampling is used.
@@
$>$ 0.5 %↑
AlphaFlow 2.87 1.63 0.53 0.66 0.85 2.64 1.55 2.29 39.0
ConfDiff 3.43 2.21 0.59 0.67 0.85 2.75 1.41 2.27 35.4
MDGen* 1.38 0.78 0.41 0.49 0.74 2.81 1.95 2.38 13.4
ConfRover-traj 3.21 1.74 0.47 0.61 0.85 2.83 1.41 2.30 41.5
ConfRover 3.63 2.23 0.51 0.64 0.85 2.62 1.39 2.28 36.6
@@
AlphaFlow 0.62 0.41 0.69 0.35 0.0 22.2
ConfDiff 0.63 0.39 0.65 0.33 0.5 6.5
MDGen* 0.50 0.28 0.57 0.26 0.3 29.0
ConfRover-traj 0.53 0.36 0.60 0.27 0.4 12.2
ConfRover 0.61 0.38 0.67 0.32 0.5 19.2
@@ E.4 Conformation Interpolation
In conformation interpolation experiment, we selected trajectories from multi-start. These trajectories exhibit sufficient conformation changes (e.g., RMSD between the start and end frames $>4$) and clear interpolation path in the PCA space. The list of selected cases are in Table 10.
@@
E.5 Retraining MDGen for Experiment-Specific Setups
Due to its non-autoregressive design and training on fixed-length trajectories, MDGen cannot directly generate sequences of varying lengths. Therefore, in our evaluation, we generate trajectories using the original settings (named MDGen-SstrideFlength) and subsample them to match the evaluation setup. However, this post-processing may introduce artifacts. To address this concern, we retrain MDGen under the evaluation settings and compare the results on the Multi-start benchmark (stride = 256) and 100 ns long trajectory simulations, as shown below in Table 11, Table 12, Table 13 and Figure 17. As shown in these experiments, no significant difference of performance for key metrics observed comparing MDGen with post-processed results and models specifically trained at the evaluation settings, suggesting no evident decrease of trajectory quality from the subsampling post-process.
@@ E.5 Retraining MDGen for Experiment-Specific Setups
Traj. Frame $\Delta$Frame Traj. Frame $\Delta$Frame

### v2

@@ Appendix B Diffusion Models on SE(3) Space
Diffusion Probabilistic Models (DPM) model complex distributions through iterative denoising. In the context of protein conformations, DPMs defined over $\mathrm{SE(3)}$ translation-rotation space have been applied for protein backbone structural generation [yimSEDiffusionModel2023, wang2024proteinconfdiff]. Following Section 2.2, $\mathbf{x}_{0}=(\mathbf{T}_{0},\mathbf{R}_{0})\in\mathrm{SE(3)}^{N}$ denotes the translations and rotations of backbone rigids in data. The diffusion processes defined in the translation and rotation subspace add noise to corrupt the data:
@@
The transition kernel of $\mathbf{T}$ satisfies $p_{t|0}(\mathbf{T}_{t}|\mathbf{T}_{0})=\mathcal{N}(\mathbf{T}_{t};\sqrt{\alpha_{t}}\mathbf{T}_{0},(1-\alpha_{t})\mathbf{I})$, where $\alpha_{t}=e^{-\int_{0}^{t}\beta_{s}ds}$. The transition kernel of $\mathbf{R}$ satisfies $p_{t|0}(\mathbf{R}_{t}|\mathbf{R}_{0})=\mathcal{IGSO}_{3}(\mathbf{R}_{t};\mathbf{R}_{0},t)$, where $\mathcal{IGSO}_{3}$ is the isotropic Gaussian distribution on $\mathrm{SO(3)}$ [yimSEDiffusionModel2023].
@@ D.1 Detailed Module Architectures
To encode the structural information of each conformation frame, we introduced the FrameEncoder, a pseudo-beta-carbon (Cβ) coordinate encoder similar to the InputEmbedding module from AlphaFlow [jingAlphaFoldMeetsFlow2024] (without diffusion time embedding). Specifically, this module first compute the pairwise distances between residues using Cβ coordinates. These distances are then binned, embedded into latent embedding, and further refined through triangular update blocks including triangle attention and multiplication updates [jumperHighlyAccurateProtein2021AF2]. See Algorithm 1 for the specifics. The resulting per-frame Cβ pair embedding $\mathbf{z}_{\text{Frame}}^{l}$ is concatenated with the pair representation from FoldingModule.
@@
In the Trajectory Module, we interleave layers of StructuralUpdate and TemporalUpdate to iteratively update the latent $\mathbf{s}^{l},\mathbf{z}^{l}]$, enabling the temporal reasoning across frames and structural refinement within each frame.
For the StructuralUpdate, we adopt a Pairformer block from AlphaFold 3 [ahdritzOpenFoldRetrainingAlphaFold22023], which jointly updates the single and pair embeddings of the current frame through structural reasoning. After the StructuralUpdate, the pair embedding is flattened from $[N,N,d]$ to $N\times N,d]$ and concatenated with the single embedding before being passed into the TemporalUpdate.
To model temporal dependencies between frames, we use a lightweight Llama architecture [touvron2023llama]. We transpose the input such that the temporal dimension is treated as the sequence axis for channel-wise self-attention across time. Rotary positional encoding [su2023roformerenhancedtransformerrotary] is applied to encode the temporal position for each frame. A causal attention mask is applied to restrict each frame to only attend to previous frames. After the temporal update, the latent embedding are reshaped and split back into single and pair embeddings.
@@
Figure 8: Architecture details of the Trajectory Module. Trajectory Module contains interleaving blocks of StructuralUpdate and TemporalUpdate (only one block of each is shown). StructuralUpdate leverages the Pairformer architecture from abramsonAccurateStructurePrediction2024, updating the pair embeddings with triangular updates and the single embeddings using with pair bias from the updated pair. The updated pair embeddings are flatten and concatenate with single embedding for channel-wise temporal update. The attention is applied along the temporal dimension and update each single and pair embedding channels independently. The embeddings from TemporalUpdate are split and reshape back into single and pair embeddings.
@@
Conformation generation, conditioned on the temporal signals from the Trajectory Module, is performed using the model architecture from ConfDiff [wang2024proteinconfdiff], As shown in Figure 9, following ConfDiff, the inputs to the denoising model include diffusion time $t$, pairwise distance between residue rigids, and residue indices (not shown). They are encoded and concatenated with the single and pair embeddings from FoldingModule. In addition, the single and pair embeddings from the Trajectory Module are projected back to the latent dimension $d_{\mathbf{s}}$ and $d_{\mathbf{z}}$, respectively, to modify the single and pair embeddings used by ConfDiff.
@@ D.2 Training and Inference Details
We train all ConfRover models on the trajectories from the ATLAS training set, following the train-validation-test split of previous works [jingAlphaFoldMeetsFlow2024, wang2024proteinconfdiff, jing2024generative_mdgen]. Specifically, we exclude the training proteins longer than 384 amino acid residues, leading to 1080 training proteins.
@@
D.3 Training and Inference Cost
Inference cost varies with protein size and benchmark setups. To measure the potential speedups from using ConfRover compared with classic MD simulation, we measured the wall-clock time required to generate 100 ns trajectories (80 frames) for ATLAS proteins of varying sizes and report the average inference time per size bucket. For comparison, we selected a representative protein from each bucket and estimated the time required to simulate 100 ns using OpenMM with implicit solvent. Both are performed on a single NVIDIA H100-80G GPU. As shown in the Table D.3, ConfRover provides significant speedup for 100 ns simulation, with even more pronounced acceleration for larger proteins.
Table 8: Runtime comparison (in minutes) across different protein sequence lengths. Speedup is computed as MD runtime divided by ConfRover runtime
Sequence length (0,150) [150,300) [300,450) [450,600) [600,724]
ConfRover 6.99 7.53 10.92 15.88 20.83
MD 104.54 207.92 386.69 651.29 1099.13
[…]
For our large-scale experimental benchmarks, multi-start trajectory simulation (2,700+ trajectories) took 8 hours and 30 minutes, while time-independent sampling (250 conformations per protein) took 3 hours and 20 minutes.
@@ E.1 Trajectory Simulation: multi-start
Here we also include the scatterplot of Pearson correlation in Figure 10 and additional metrics from the multi-start experiments at different strides in Table 9 (one experimental included). Across different strides, ConfRover models consistently outperforms MDGen in recovering the correct level of dynamics.
[…]
Table 9: Additional metrics from the multi-start benchmark. Results for different strides are shown in separate blocks. The better score in each block is highlighted in bold (excluding diversity). One inference run per model is used for this comparison.
@@ E.3 Time-independent Conformation Sampling
We follow the evaluation protocol of ye2024proteinbench to assess time-independent conformation sampling on the ATLAS test set. For each protein, 250 independent samples are generated. Since MDGen does not support time-independent sampling, we approximate its performance using samples from its 100-ns trajectory, serving as a sequential-sampling baseline. The performance of state-of-the-art models, AlphaFlow and ConfDiff, is taken from Table 10 of ye2024proteinbench, using their best-performing variants: AlphaFlow-MD and ConfDiff-Open-MD. Full results are shown in Table 10.
Table 10: Performance on time-independent conformation generation on ATLAS. A total of 250 conformations were sampled for each protein, and the mean and standard deviation of metrics are computed from five independent runs. The best performance is highlighted in bold, and the second-best is underlined.ConfRover-traj and MDGen are trained to exclusive generate trajectories. MDGen does not support time-independent sampling and the metrics are evaluated on sequential sampling result.
@@
>> 0.5 %↑
[…]
@@
[…]
@@ E.4 Conformation Interpolation
In conformation interpolation experiment, we selected trajectories from multi-start. These trajectories exhibit sufficient conformation changes (e.g., RMSD between the start and end frames >4>4 Å) and clear interpolation path in the PCA space. The list of selected cases are in Table 11.
@@
E.5 Extended Baseline Comparisons
E.5.1 Retraining MDGen for Experiment-Specific Setups
Due to its non-autoregressive design and training on fixed-length trajectories, MDGen cannot directly generate sequences of varying lengths. Therefore, in our evaluation, we generate trajectories using the original settings (named MDGen-S{stride}F{length}) and subsample them to match the evaluation setup. However, this post-processing may introduce artifacts. To address this concern, we retrain MDGen under the evaluation settings and compare the results on the Multi-start benchmark (stride = 256) and 100 ns long trajectory simulations, as shown below in Table 12, Table 13, Table 14 and Figure 16. Experimental results show no significant difference of performance for key metrics observed comparing MDGen with post-processed results and models specifically trained at the evaluation settings, suggesting no evident decrease of trajectory quality from the subsampling post-process. We use one inference run per model for this experiment.
@@ E.5.1 Retraining MDGen for Experiment-Specific Setups
$\DeltaFrame Traj. Frame Δ\DeltaFrame$
@@
E.5.2 Internal Reproduction of AlphaFolding
Available baseline models for protein trajectory generation are limited at the time of this work, we attempted to reproduce AlphaFolding [cheng20244d4ddiff], a diffusion-based model that generate short trajectories (i.e., blocks) and can be extended to longer through iterative generation. We followed the authors’ official implementation  [^2] https://github.com/fudan-generative-vision/dynamicPDB/tree/main/applications/4d_diffusion, setting the motion token count to 2 and the generation horizon (block length) to 16 frames. To improve generation efficiency, we increased the stride from the default 1 to 40, matching the setup in MDGen. Training AlphaFolding on the full ATLAS dataset resulted in out-of-memory error on NVIDIA A100-80GB GPU. Therefore, we adopted the authors’ filtering criterion of a maximum sequence length of 256 residues. Similarly, we encountered the out-of-memory error for the five largest proteins during inference. The model was trained for 65K steps where we saw convergence. For evaluation, we iteratively extended the 16-frame outputs to generate 256 frames and retained the first 250 frames to match the 100 ns simulation setting.
We compared AlphaFolding with MDGen and ConfRover on the ATLAS 100 ns simulation task, using one inference run per model. As shown in Table 15, both AlphaFolding and ConfRover outperform MDGen in capturing the principal coordinates of the dominant dynamics, although AlphaFolding slightly lags behind ConfRover. However, as shown in Table 16, AlphaFolding produces lower-quality conformations, exhibiting inflated backbone and side-chain rotamer outliers as well as higher clash rates. We found more evident degradations from error accumulation (Table 17), likely due to its iterative block-wise extension, which conditions only on the last frame of the preceding block. In contrast, MDGen employs non-autoregressive attention across all frames, while ConfRover maintains a full attention history via KV cache. Owing to the noisy nature of its generated trajectories, AlphaFolding also shows inflated recall in conformational state recovery (Table 18), as the increased structural noise likely leads to artificially higher coverage in state space.
Although our implementation is a reproduction of [cheng20244d4ddiff] and may not faithfully reflect the authors’ original model, this analysis suggests the potential limitations of block-extension-based trajectory models including AlphaFolding.
Table 15: Pearson correlations of principal dynamic modes (PC1 and PC2) between sampled and reference trajectories, evaluated at varying lag times ($\Delta t$, in frames). The best scores are highlighted in bold.
$\Delta t=1$ $\Delta t=5$ $\Delta t=10$ $\Delta t=20$
PC 1 MDGen 0.11 0.12 0.10 0.12
PC 1 AlphaFolding 0.15 0.15 0.16 0.18
PC 1 ConfRover 0.19 0.17 0.19 0.19
PC 2 MDGen 0.12 0.10 0.10 0.11
PC 2 AlphaFolding 0.18 0.15 0.16 0.20
PC 2 ConfRover 0.19 0.17 0.18 0.17
Table 16: Conformation geometric quality evaluated using MolProbity. Un-relaxed conformations are used for evaluation. The best scores are highlighted in bold.
Ramachandran
outliers % ($\downarrow$)
Rotamer
outliers % ($\downarrow$)
Clash
score ($\downarrow$)
RMS
bonds ($\downarrow$)
RMS
angles ($\downarrow$)
MolProbity
score ($\downarrow$)
AlphaFolding 2.91 15.37 151.35 0.07 4.76 3.94
MDGen 1.87 2.85 128.08 0.04 3.14 3.28
ConfRover 1.75 3.30 76.89 0.05 3.69 3.00
Table 17: Average Ramachandran outliers across trajectory frame ranges. Compared with MDGen and ConfRover, AlphaFolding exhibits an increased level of backbone outliers when simulating longer trajectories.
Frame range (0,15] (15,31] (31,47] (47,63] (63,79]
AlphaFolding 1.77 2.78 3.21 3.34 3.57
MDGen 1.59 1.85 1.96 1.95 2.05
ConfRover 1.34 1.61 1.77 1.93 2.08
Table 18: Recovery of conformational states in the ATLAS 100 ns simulation experiment. Although AlphaFolding shows the highest coverage, this may due to lower sample quality that artificially inflates diversity.
JS-Dist $\downarrow$ Recall $\uparrow$ F1 $\uparrow$
AlphaFolding 0.47 0.51 0.65
MDGen 0.55 0.29 0.43
ConfRover 0.51 0.42 0.58
E.6 Extension to Masked Sequence Modeling
While the causal transformers used in ConfRover enables efficient autoregressive training and generation, proposed framework can be extended to other sequence modeling paradigm, such as masked sequence modeling. Similar to li2024autoregressivemar, sequence models trained with bidirectional attention and masked sequence modeling can perform autoregressive generation by iteratively predicting frames in a specified order. At each iteration, the generated frame replaces its corresponding mask and joins the input sequence for the next iteration. We trained a variant ConfRover-mask following this approach. Specifically, we replaced causal attention with bidirectional attention in the transformer and implemented a scheduled masking strategy for training, linearly increasing the mask rate from 45% to 88%.
[…]
MDGen 0.55 0.45 0.40
ConfRover 0.77 0.63 0.53
ConfRover-mask 0.71 0.60 0.49
PCA 2D
[…]
MDGen 0.16 0.11 0.10
ConfRover 0.75 0.50 0.43
ConfRover-mask 0.73 0.49 0.41